OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 1) and (3, 4)

OpenStudy (anonymous):

You need to find the slope first which is rise over run. (y2-y1)/(x2-x1)

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

Then you sub that into the formula for y = mx + b. where m is the slope that you found, and you sub in a point on the line. You know that (1,1) is a point on the line

OpenStudy (anonymous):

So solve for b :) 1 = (2/3) 1 + b Do you understand how I got that equation?

OpenStudy (anonymous):

I think I got iit

OpenStudy (anonymous):

Whoops, it's (3/2) for the slope

OpenStudy (anonymous):

You said 2/3 but remember rise over run

OpenStudy (anonymous):

is 1 = (2/3) 1 + b written in the standard form?

OpenStudy (anonymous):

No, you have to solve for b. Also the slope you're using is not correct. Remember, rise over run

OpenStudy (anonymous):

The standard form is y = mx + b where you know m (the slope) and b (the y intercept).